# The Dielectric Response of Plasmas with Arbitrary Gyrotropic Velocity Distributions

### v1

@@ Abstract
Hot and tenuous plasmas are frequently far from local thermodynamic equilibrium, necessitating sophisticated methods for determining the associated plasma dielectric tensor and normal mode response. The Arbitrary Linear Plasma Solver (ALPS) is a numerical tool for calculating such responses of plasmas with arbitrary gyrotropic background velocity distribution functions (VDFs). In order to model weakly and moderately damped plasma waves accurately, we have updated to the code to use an improved analytic continuation enabled by a polynomial basis representation. We demonstrate the continuity of solutions to the linear Vlasov–Maxwell dispersion relation between bi-Maxwellian and arbitrary VDF representations and evaluate the influence of VDF structure on mode polarization and wave power emission and absorption.
@@ I Introduction
The initial ALPS implementation is described by Verscharen et al. (2018). Recent additions to the code enable a more accurate representation of moderately damped solutions for VDFs that deviate from a bi-Maxwellian background. The additions also include the calculation of the associated eigenfunctions and a partitioning of the species heating rates into the component mechanisms, e.g. Landau, Transit Time, and Cyclotron damping. We explore the impact of the updated representation for the calculation of analytic continuation of the Landau contour integral on weakly and moderately damped modes in § II. We demonstrate the continuity of solutions between bi-Maxwellian fits and the ALPS direct calculation, as well as the changes for the underlying eigenfunctions and damping rates, in § III. Details of the underlying numerics are found in Appendices A, B, and C.
II Improved Hybrid Analytic Continuation
@@
When the solutions are damped ($\gamma<0$, where $\gamma=\mathbb{I}\mathrm{m}(\omega)$) the integration of Eqn. 2 necessitates an analytic continuation of $f_{0,s}$ into the complex $p_{\parallel}$ plane. This is straightforward if $f_{0,s}$ is a known analytical function which can be evaluated at a complex $p_{\parallel}$. [^1] The relativistic case is more complex, given the non-trivial momentum dependence of the resonant denominator in that limit; see §3.3 by Verscharen et al. (2018) for details, where we use the method suggested by Lerche:1967 to transform from $(p_{\parallel},p_{\perp})$ to $(\Gamma,\bar{p}_{\parallel})$ and make the solution tractable. For the non-analytical case, as we only have values for $f_{0,s}$ for $\mathbf{p}\in\mathbb{R}$, we use two complementary schemes to extend the distribution function numerically. The initial implementation of ALPS is a fit of $f_{0,s}$ to a (typically small) number of analytical functions, § II.1. In this work, we describe a newly implemented polynomial basis representation using a generalized linear least squares approach, § II.2. In both cases, we follow the prescription in Landau (1946) for the integration of a function $G$ over a contour $C_{L}$ that lies below the complex poles of the integrand:
@@
=$=\begin{cases}\int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})&\text{if }\gamma>0,\\ \mathcal{P}\int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})+i\pi\sum\mathrm{Res}_{A}(G)&\text{if }\gamma=0,\\ \int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})+2i\pi\sum\mathrm{Res}_{A}(G)&\text{if }\gamma<0,\end{cases}$
@@ II.1 Fit Function Representation
However, when applying the code to more complicated VDFs, two recurrent difficulties can arise. First, even when a good set of fit initializations were selected, the fitting routine could fail at larger $p_{\perp}$ values, typically due to pathological but not unphysical variations of the VDF as a function of $p_{\perp}$, e.g. the disappearance of a secondary population or the appearance of an additional population at larger $p_{\perp}$. This failure would cause the entire fitting procedure to halt. Second, even when the fits converge for all $p_{\perp}$, the small number of functions used for the analytical representation lead to inaccuracies in the representation of the VDF, producing discrepancies as dispersion relations cross from $\gamma>0$ to $\gamma<0$ or vice versa.
These discrepancies arise for two reasons. First, the momentum derivatives of $f_{0,s}$ and the fit function can differ when evaluated at the resonant velocity $v_{\textrm{res}}=(\omega-n\Omega_{s})/k_{\parallel}$. Secondly, the total velocity moments of the fit function can differ from the velocity moments of the actual VDF, leading to an effective net charge and current that is different for $\gamma>0$ and $\gamma<0$. In addition to producing incorrect mode structure for damped solutions, these discrepancies can lead to a sharp edge at $\gamma=0$ that causes the numerical Newton-secant search used to follow dispersion relations with varying wavevector to jump to a different solution or lose the physical solutions entirely due to the large gradient in $\Lambda$. Such a jump is illustrated in the surfaces associated with low-order fits in Fig. 1, e.g. the grey line in the lower right panel. These issues have led us to the development of a new approach to perform the hybrid analytic continuation, a polynomial basis representation.
II.2 Polynomial Basis Representation
@@
As $T_{n}$ are bound between $[-1,1]$, we perform a change of variables from $p_{\parallel}$ to $x$ through
@@
The bounded nature of $T_{n}$ is ideally suited to the limited range of $p_{\parallel}$ used as inputs for ALPS. Chebyshev polynomial representation performs better than other classes of orthogonal polynomials, e.g. the Hermite or Legendre polynomials, not shown. We model $\log_{10}[f_{0,s}]$ rather than $f_{0,s}$ to capture the full range of the VDF amplitude and not solely the structure near the peak of the distribution. We emphasize that, as with the functional fitting approach discussed in § II.1, this is a purely one-dimensional method, representing $f_{0,s}(p_{\parallel})$ separately for the $n_{\perp}$ rows in the input VDF grid; we write the VDF slice that we represent as $y_{i}$, which is known at $N$ values of $x$.
@@
If the errors in the knowledge of $y_{i}$ are known, one can set $\sigma_{i}$ to non-unity values; in the current implementation, $\sigma_{i}=1\forall x_{i}$. Eqn. 18 is solved using the dgemm, dgemv, and dgesv routines from the LAPACK library.
@@
The polynomial basis method produces superior results compared to the functional fitting method for highly structured VDFs, both in terms of the smoothness of $\Lambda(\omega)$ across $\gamma=0$ due to improved accuracy of $\partial_{v}f_{0,j}$ for all $\omega$ and $\bm{k}$ values, as well as in correctly capturing the total density and current of the VDFs. To illustrate this, Fig. 1 shows the representation of a slice of $f_{0,p}(p_{\parallel})$ for $p_{\perp}=0$ using the GLLS method for orders $N=(10,20,30,40,50,60)$ and the best fit to the sum of two bi-Maxwellians. Also shown are the relative difference between the fit or GLLS representation and the actual VDF slide. We consider a distribution similar to one from Walters et al. (2023), interpolated from Wind SWE observations to a fixed grid of $300\times 600$ points in $(p_{\perp}\times p_{\parallel})$. Looking at the $p_{\perp}=0$ slice, neither the bi-Maxwellian fit (purple) nor the lowest order GLLS representation (grey) of $f_{p}(p_{\parallel})$ match the actual distribution, with the relative error approaching or exceeding order unity. Moving to higher order representations, the relative error decreases by several orders of magnitude; the highest order we consider, $\mathcal{O}(60)$, is well below the number of resolved points in $p_{\parallel}$, ensuring we are not over-fitting the distribution.
In the right-hand panels of Fig. 1, we plot isocontours of the magnitude of $\log_{10}[\Lambda(\omega_{\textrm{r}},\gamma)]$, as well as the associated solutions satisfying $\Lambda=0$, for a wavevector of $(k_{\perp},k_{\parallel})d_{p}=(0.001,0.0025)$. The isocontours for both the lowest, $\mathcal{O}(10)$, and highest, $\mathcal{O}(60)$, representations are shown. For the $\mathcal{O}(10)$ case (grey dashed lines), there is a stark discontinuity across $\gamma=0$, indicating that the representation is not sufficiently accurate for the analytic continuation to enable a smooth transition from unstable to stable solutions. For the $\mathcal{O}(60)$ case (red), the isocontours are continuous across $\gamma=0$; the transition between these two orders over a limited region of complex frequency is illustrated in the upper right inset, illustrating the convergence towards a smooth surface with higher polynomial order.
@@
III Comparing Dispersion Relations
III.1 Validating against bi-Maxwellian Models
@@
where the left handed waves (forward Alfvén and backward fast) have $n=1$ and the right handed waves (backwards Alfvén and forward fast) have $n=-1$, are effectively identical between the solutions from the two models.
@@
What drives these differences can be identified using quasilinear theoryKennel and Petschek (1966); Verscharen and Chandran (2013). The resonant coupling of the wave to the VDF can be represented in terms of the operator
@@
applied to $f_{p}(v_{\perp},v_{\parallel})$ To illustrate the structures of the VDF responsible for growth and damping, we construct a $k_{\parallel}$-dependent function, applying $\mathcal{G}$ to $f_{p}$ at $v_{\perp}$ and the resonant parallel velocity associated with each wavevector,
@@
Velocities are normalized to the Alfvén velocity $v_{A}=B/\sqrt{4\pi n_{p}m_{p}}$. This function is illustrated in the bottom two rows of Fig. 3. As this calculation is only valid when $|\gamma|/|\omega_{\textrm{r}}|<1/e$, and only useful when the resonant velocity samples the input VDF, we restrict ourselves to the limited range or wavevectors satisfying both these conditions. The sign of this function at each velocity value indicates whether that point in phase space emits or absorbs power in interaction with the given wave. As illustrated by Walters et al. (2023), the sign of the function integrated over $v_{\perp}$ indicates whether the species has a net absorption or emission at that parallel scale.
@@ IV Continuity of Solutions
Figure 6: (Left) Input proton VDFs used for the ALPS calculations, including the Observed Wind VDF $f_{Obs.}$, best-fit two-bi-Maxwellian $f_{C\&B}$, and one intermediate model. (Right) Real part of the frequency $\omega_{\textrm{r}}/\Omega_{p}$, damping (dashed lines) and growth (solid) rates $\gamma/\omega_{\textrm{r}}$, the density-$B_{z}$ correlation, electric field polarization $\mathbf{P}_{xy}$, and $v_{\textrm{res}}$ as a function of $k_{\parallel}d_{p}$ for fixed $k_{\perp}d_{p}$ for the forward and backward Alfvén solutions using the continuous set of VDFs.
@@
Figure 7: Dispersion relations for the forward propagating fast and slow magnetoacoustic modes for the core-and-beam VDF $f_{C\&B}$, the observed VDF $f_{Obs.}$, and intermediate VDF representations $f_{i}$. (left panel) Parametric curves of $[\omega_{\text{r}},\gamma](k_{\parallel},d_{p})$ show distinct connections between the MHD-scale and ion-scale solutions for the observed and core-and-beam models. (right) Variation in $C(\delta n_{p},\delta B_{\parallel})$ of the fast and slow solutions as a function of parallel wavevector.
@@ Acknowledgments
The authors thank the UCL Open Source Software Sustainability Project for their assistance in publishing ALPS as an open source code. This study benefited from support by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 24-612 ("Excitation and dissipation of kinetic-scale fluctuations in space plasmas"). K.G.K was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) and NASA grant No. 80NSSC19K091 and 80NSSC240171. D.V. is supported by STFC Consolidated Grant ST/W001004/1.
@@ Appendix A Using bi-Maxwellian and Cold Plasma Models for Select Species
[…]

### v2

@@ Abstract
Hot and tenuous plasmas are frequently far from local thermodynamic equilibrium, necessitating sophisticated methods for determining the associated plasma dielectric tensor and normal mode response. The Arbitrary Linear Plasma Solver (ALPS) is a numerical tool for calculating such responses of plasmas with arbitrary gyrotropic background velocity distribution functions (VDFs). In order to model weakly and moderately damped plasma waves accurately, we have updated the code to use an improved analytic continuation enabled by a polynomial basis representation. We demonstrate the continuity of solutions to the linear Vlasov–Maxwell dispersion relation between bi-Maxwellian and arbitrary VDF representations and evaluate the influence of VDF structure on mode polarization and wave power emission and absorption.
@@ I Introduction
The initial ALPS implementation is described by Verscharen et al. (2018). Recent additions to the code include the calculation of the associated eigenfunctions and a partitioning of the species heating rates into the component mechanisms, e.g. Landau, Transit Time, and Cyclotron damping, as well as the option to treat selected plasma components as bi-Maxwellians for numerical expediency. These changes are discussed in § II. The original code struggled with the evaluation of the Landau contour integral essential for characterizing damped solutions for some VDFs with strong non-Maxwellian features. We present updates that enable a more accurate representation in this evaluation for moderately damped solutions for such VDFs. We explore the impact of the updated representation on weakly and moderately damped modes in § III. We demonstrate the continuity of solutions between bi-Maxwellian fits and the ALPS direct calculation, as well as the changes for the underlying eigenfunctions and damping rates, in § IV. Details of the underlying numerics are found in Appendices A, B, and C.
II Dispersion Relations for Arbitrary Distributions
@@
III Improved Hybrid Analytic Continuation
When the solutions are damped ($\gamma<0$, where $\gamma=\mathbb{I}\mathrm{m}(\omega)$) the integration of Eqn. 2 necessitates an analytic continuation of $f_{0,s}$ into the complex $p_{\parallel}$ plane. This is straightforward if $f_{0,s}$ is a known analytical function which can be evaluated at a complex $p_{\parallel}$. [^1] The relativistic case is more complex, given the non-trivial momentum dependence of the resonant denominator in that limit; see §3.3 by Verscharen et al. (2018) for details, where we use the method suggested by Lerche:1967 to transform from $(p_{\parallel},p_{\perp})$ to $(\Gamma,\bar{p}_{\parallel})$ and make the solution tractable. For the non-analytical case, as we only have values for $f_{0,s}$ for $\mathbf{p}\in\mathbb{R}$, we use two complementary schemes to extend the distribution function numerically. The initial implementation of ALPS is a fit of $f_{0,s}$ to a (typically small) number of analytical functions, § III.1. In this work, we describe a newly implemented polynomial basis representation using a generalized linear least squares approach, § III.2. In both cases, we follow the prescription in Landau (1946) for the integration of a function $G$ over a contour $C_{L}$ that lies below the complex poles of the integrand:
@@ III Improved Hybrid Analytic Continuation
$=\begin{cases}\int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})&\text{if }\gamma>0,\\ \mathcal{P}\int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})+i\pi\sum\mathrm{Res}_{A}(G)&\text{if }\gamma=0,\\ \int_{-\infty}^{+\infty}\mathrm{d}p_{\parallel}G(p_{\perp},p_{\parallel})+2i\pi\sum\mathrm{Res}_{A}(G)&\text{if }\gamma<0,\end{cases}$
@@ III.1 Fit Function Representation
However, when applying the code to more complicated VDFs, two recurrent difficulties can arise that necessitate the development of a new representation method. First, even when a good set of fit initializations were selected, the fitting routine could fail at larger $p_{\perp}$ values, typically due to pathological but not unphysical variations of the VDF as a function of $p_{\perp}$, e.g. the disappearance of a secondary population or the appearance of an additional population at larger $p_{\perp}$. This failure would cause the entire fitting procedure to halt. Second, even when the fits converge for all $p_{\perp}$, the small number of functions used for the analytical representation lead to inaccuracies in the representation of the VDF, producing discrepancies as dispersion relations cross from $\gamma>0$ to $\gamma<0$ or vice versa.
These discrepancies arise for two reasons. First, the momentum derivatives of $f_{0,s}$ and the fit function can differ when evaluated at the resonant velocity $v_{\textrm{res}}=(\omega-n\Omega_{s})/k_{\parallel}$. Secondly, the total velocity moments of the fit function can differ from the velocity moments of the actual VDF, leading to an effective net charge and current that is different for $\gamma>0$ and $\gamma<0$. In addition to producing incorrect mode structure for damped solutions, these discrepancies can lead to a sharp edge at $\gamma=0$ that causes the numerical Newton-secant search used to follow dispersion relations with varying wavevector to jump to a different solution or lose the physical solutions entirely due to the large unphysical gradient in $\Lambda$. Such a jump is illustrated in the surfaces associated with low-order fits in Fig. 1, e.g. the grey line in the lower right panel.
These issues with the fit function representation for VDFs that are significantly different from a finite superposition of Maxwellian functions in the $p_{\parallel}$ dimension have led us to the development of a new approach to perform the hybrid analytic continuation, combining numerical integration of the input VDF with a polynomial basis representation when the dispersion relation requires evaluation of the VDF at complex momentum values.
III.2 Polynomial Basis Representation
@@
As $T_{n}$ are bound between [1,1][-1,1], we perform a change of variables from $p_{\parallel}$ to $x$ through
@@
The bounded nature of $T_{n}$ is ideally suited to the limited range of $p_{\parallel}$ used as inputs for ALPS. Chebyshev polynomial representation performs better than other classes of orthogonal polynomials, e.g. the Hermite or Legendre polynomials, not shown. We model $\log_{10}[f_{0,s}]$ rather than $f_{0,s}$ to capture the full range of the VDF amplitude and not solely the structure near the peak of the distribution. We emphasize that, as with the functional fitting approach discussed in § III.1, this is a purely one-dimensional method, representing $f_{0,s}(p_{\parallel})$ separately for the $n_{\perp}$ rows in the input VDF grid; we write the VDF slice that we represent as $y_{i}$, which is known at $N$ values of $x$.
@@
If the errors in the knowledge of $y_{i}$ are known, one can set $\sigma_{i}$ to non-unity values; in the current implementation, $\sigma_{i}=1\forall x_{i}$. Eqn. 18 is solved using the dgemm, dgemv, and dgesv routines from the LAPACK library. As the GLLS solution is calculated once for an input VDF, there is no need to implement more efficient existing schemes for determining $a_{kj}$ that exploit connections between Fourier and Chebyshev representations. (e.g. Mason and Handscomb, 2002)
@@
The polynomial basis method produces superior results compared to the functional fitting method for highly structured VDFs, both in terms of the smoothness of $\Lambda(\omega)$ across $\gamma=0$ due to improved accuracy of $\partial_{p}f_{0,j}$ for all $\omega$ and $\bm{k}$ values, as well as in correctly capturing the total density and current of the VDFs. To illustrate this, Fig. 1 shows the representation of a slice of $f_{0,p}(p_{\parallel})$ for $p_{\perp}=0$ using the GLLS method for orders $N=(10,20,30,40,50,60)$ and the best fit to the sum of two bi-Maxwellians. Also shown are the relative difference between the fit or GLLS representation and the actual VDF slide. We consider a distribution similar to one from Walters et al. (2023), interpolated from Wind SWE observations to a fixed grid of $300\times 600$ points in $(p_{\perp}\times p_{\parallel})$. Looking at the $p_{\perp}=0$ slice, neither the bi-Maxwellian fit (purple) nor the lowest order GLLS representation (grey) of $f_{p}(p_{\parallel})$ match the actual distribution, with the relative error approaching or exceeding order unity. Moving to higher order representations, the relative error decreases by several orders of magnitude; the highest order we consider, $\mathcal{O}(60)$, is well below the number of resolved points in $p_{\parallel}$, ensuring we are not over-fitting the distribution.
In the right-hand panels of Fig. 1, we plot isocontours of the magnitude of $\log_{10}[\Lambda(\omega_{\textrm{r}},\gamma)]$, as well as the associated solutions satisfying $\Lambda=0$, for a wavevector of $(k_{\perp},k_{\parallel})d_{p}=(0.001,0.0025)$. The isocontours for both the lowest, $\mathcal{O}(10)$, and highest, $\mathcal{O}(60)$, representations are shown. For the $\mathcal{O}(10)$ case (grey dashed lines), there is a stark discontinuity across $\gamma=0$, indicating that the VDF representation is not sufficiently accurate in its representation of the bulk velocity moments and resonant momentum gradients involved in the principal-value integration, Eqn. 7, to enable a smooth transition between growing and damping solutions. For the $\mathcal{O}(60)$ case (red), the isocontours are continuous across $\gamma=0$; the transition between these two orders over a limited region of complex frequency is illustrated in the upper right inset, illustrating the convergence towards a smooth surface with higher polynomial order.
@@
IV Comparing Dispersion Relations
IV.1 Validating against bi-Maxwellian Models
@@
Figure 3: Input proton VDFs used for the ALPS calculations, including the best-fit two-bi-Maxwellian $f_{C\&B}$ (top), observed Wind VDF $f_{Obs.}$ (bottom), and one intermediate model, Eqn. 25 (middle).
@@
where the left handed waves (forward Alfvén and backward fast) have $n=1$ and the right handed waves (backwards Alfvén and forward fast) have n=1n=-1, are effectively identical between the solutions from the two models.
@@
What drives these differences can be identified using quasilinear theoryKennel and Engelmann (1966); Verscharen and Chandran (2013). The resonant coupling of the wave to the VDF can be represented in terms of the operator
@@
applied to $f_{p}(v_{\perp},v_{\parallel})$. To illustrate the structures of the VDF responsible for growth and damping, we construct a $k_{\parallel}$-dependent function, applying $\mathcal{G}$ to $f_{p}$ at $v_{\perp}$ and the resonant parallel velocity associated with each wavevector,
@@
Velocities are normalized to the Alfvén velocity $v_{A}=B/\sqrt{4\pi n_{p}m_{p}}$. This function is illustrated in the bottom two rows of Fig. 4. As this calculation is only valid in the weak damping or growth limitKennel and Engelmann (1966) $|\gamma|/|\omega_{\textrm{r}}|<1/e$, and only useful when the resonant velocity samples the input VDF, we restrict ourselves to the limited range or wavevectors satisfying both these conditions. The sign of this function at each velocity value indicates whether that point in phase space emits or absorbs power in interaction with the given wave. As illustrated by Walters et al. (2023), the sign of the function integrated over $v_{\perp}$ indicates whether the species has a net absorption or emission at that parallel scale.
@@ V Continuity of Solutions
Figure 7: Real part of the frequency $\omega_{\textrm{r}}/\Omega_{p}$, damping (dashed lines) and growth (solid) rates $\gamma/\omega_{\textrm{r}}$, the density-$B_{z}$ correlation, electric field polarization $\mathbf{P}_{xy}$, and $v_{\textrm{res}}$ as a function of $k_{\parallel}d_{p}$ for fixed $k_{\perp}d_{p}$ for the forward and backward Alfvén solutions using the continuous set of VDFs.
@@
Figure 8: Dispersion relations for the forward propagating fast and slow magnetoacoustic modes for the core-and-beam VDF $f_{C\&B}$, the observed VDF $f_{Obs.}$, and intermediate VDF representations $f_{i}$. (left panel) Parametric curves of $\omega_{\text{r}},\gamma](k_{\parallel},d_{p})$ show distinct connections between the MHD-scale and ion-scale solutions for the observed and core-and-beam models. (right) Variation in $C(\delta n_{p},\delta B_{\parallel})$ of the fast and slow solutions as a function of parallel wavevector.
@@ Acknowledgments
The authors thank the UCL Open Source Software Sustainability Project for their assistance in publishing ALPS as an open source code. This study benefited from support by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 24-612 ("Excitation and dissipation of kinetic-scale fluctuations in space plasmas"). K.G.K was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) and NASA grant No. 80NSSC19K0912 and 80NSSC24K0171. D.V. is supported by STFC Consolidated Grant ST/W001004/1.
@@ Appendix A Using bi-Maxwellian and Cold Plasma Models for Select Species
$Z_{0}(\zeta)=\frac{1}{\sqrt{\pi}}\mathcal{P}\int\limits_{-\infty}^{\infty}dz\,\frac{e^{-z^{2}}}{z-\zeta}+i\sqrt{\pi}\,\text{sgn}(k_{\parallel})e^{-\zeta^{2}},$ (31)